AndreaRivera:

The product of sin 30° and sin 60° is same as the product of ______and _______.

687vhlovdjfl:

options?

Imagine:

90* 120*?

jhonyy9:

so sin 30° = 1/2 and sin 60° = sqrt2 /3 and hope you know again that cos 60° = 1/2 and cos 30° = sqrt2 /3 hope this will help you